### 04 Mar 2026 v1

# Liouville phenomenon for the Klein-Gordon equation in $1+1$ dimensions

Abstract.
We study the Klein-Gordon equation in one spatial and one temporal dimension. Physically, this equation describes the wave function of a relativistic spinless boson with positive rest mass. Mathematically, this is the most elementary hyperbolic partial differential equation, after the wave equation itself. Relative to the point at the origin, the spacetime splits according to the light cones, and we find four quarter-planes, two of which are timelike while the remaining two are spacelike. Not unexpectedly, the solutions behave quite differently in the two types of quarter-planes. It turns out that the spacelike quarterplanes exhibit a Liouville phenomenon, where insufficient growth forces the solution to display a certain kind of symmetry, where the values along the two linear edges are in a one-to-one relation. This phenomenon shares features with the classical Liouville theorem in the theory of entire functions as well as the Phragmén-Lindelöf principle for harmonic functions.
Key words and phrases:
Liouville theorem, Klein-Gordon equation, Phragmén-Lindelöf principle, hyperbolic equation
1991 Mathematics Subject Classification:
Primary 35L10, 35Q40, 35B53
1. Introduction
1.1. Classical backgrund
The Klein-Gordon equation describes the wavefunction of a relativistic spinless boson with positive rest mass. In one spacial and one temporal dimension, it is advantageous to swith to using the characteristic directions to define alternative coaordinates, which after rescaling gives us the equation
(1.1) $u^{\prime\prime}_{xy}+u=0,$
where $u(x,y)$ is the wavefunction in question. The characteristic directions are now the lines parallel to the $x$-axis as well as those parallel to the $y$-axis. In this context, the corresponding wave equation reads $u^{\prime\prime}_{xy}=0$ with general solution of the form $u(x,y)=f(x)+g(y)$. Here, in principle, the solution exhibits no extra smoothness and if we look for distributional solutions, then $f,g$ should be thought of as arbitrary distributions of one variable. We shall be concerned mainly with continuous solutions $u$ to the wave equation as well as to the Klein-Gordon equation. In the context of the wave equation with general solution $u(x,y)=f(x)+g(y)$, the univariate functions $f,g$ are then continuous as well. The Lorentz group describes the linear transformations of the $(x,y)$-plane which preserve the shape of the Klein-Gordon equation. In our simplified $1+1$ dimensional context, it consists of the $2\times 2$ matrices
$A=\begin{pmatrix}\lambda&0\\ 0&\lambda^{-1}\end{pmatrix}\quad\text{or}\quad\begin{pmatrix}0&\lambda\\ \lambda^{-1}&0\end{pmatrix}$,
where $\lambda\in\mathbb{R}$, $\lambda\neq 0$. In addition to such linear Lorentz transformations, the Klein-Gordon equation is also invariant under all translations $(x,y)\mapsto(x+a,y+b)$, for $(a,b)\in\mathbb{R}^{2}$, which entails that the equation is invariant under the action of the semidirect sum of the translation group and the Lorentz group, called the Poincaré group.
It is natural to consider the Darboux-Goursat problem for such hyperbolic equations. For the Klein-Gordon equation, the Darboux-Goursat problem associated with the given characteristic axes reads
(1.2) $\begin{cases}u^{\prime\prime}_{xy}+u=0,\\ u(x,0)=f(x),\quad u(0,y)=g(y).\end{cases}$
Here, given that we ask for a continuous solution $u$, we should require that $f,g$ be continuous with $f(0)=g(0)$. As a side note, we should mention that a discretized version of the Darboux-Goursat problem (1.2) was considered recently in [3], for a restricted class of bounded solutions $u$. Equations of the type (1.2) were of interest to Bernhard Riemann, and for this reason the solution is associated with his name. Before we come to that, we need to introduce certain Bessel-type functions. For $a=0,1,2,\ldots$, the bivariate Bessel function
$J_{0,a}(y,x):=J_{a,0}(x,y):=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{(j+a)!j!}\,x^{j+a}y^{j}=\bigg(\frac{x}{y}\bigg)^{a/2}J_{a}(2\sqrt{xy})$
is an entire function of two variables, where on the right-hand side the function $J_{a}$ denotes the usual Bessel function. We quickly verify that
$\partial_{x}\partial_{y}J_{a,0}(x,y)+J_{a,0}(x,y)=0,\qquad\text{if}\,\,\,a=0\,\,\,\text{or}\,\,\,b=0,$
so that these functions constitute solutions to the Klein-Gordon equation. Moreover, since we have agreed that $J_{a,0}(x,y)=J_{0,a}(y,x)$, the function $J_{0,a}(x,y)$ solves the Klein-Gordon equation as well. Note that differentiation has the effect
$\partial_{x}J_{a,0}(x,y)=J_{a-1,0}(x,y),\qquad a\geq 1,$
while
$\partial_{x}J_{0,0}(x,y)=-J_{0,1}(x,y)$
and, similarly, we have
$\partial_{y}J_{a,0}(x,y)=-J_{a+1,0}(x,y),\qquad a\in\mathbb{Z}_{\geq 0}.$
Finally, we note that restriction to the lines $x=0$ and $y=0$ gives
$J_{0,0}(0,y)=1,\quad J_{a,0}(0,y)=0,\,\,\,\text{if}\,\,\,a\geq 1,$
and
$J_{a,0}(x,0)=\frac{x^{a}}{a!}.$
Definition 1.1.
Riemann’s solution to the Darboux-Goursat problem is given by the expression
(1.3) $\mathbf{R}[f,g](x,y):=f(x)+g(y)-f(0)J_{0,0}(x,y)-\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s\\ -\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t.$
We now make some observations which are classical. To begin with, we note that
(1.4) $\mathbf{R}[f,g](x,0)=f(x)+g(0)-f(0)J_{0,0}(x,0)-\int_{0}^{x}J_{0,1}(x-s,0)f(s)\,\mathrm{d}s=f(x)$
since $f(0)=g(0)$, $J_{0,0}(x,0)=1$, and $J_{1,0}(0,x-s)=0$. By interchanging the roles of $x$ and $y$ in the calculation, we find likewise that
(1.5) $\mathbf{R}[f,g](0,y)=f(0)+g(y)-f(0)J_{0,0}(0,y)-\int_{0}^{y}J_{1,0}(0,y-t)g(t)\,\mathrm{d}t=g(y).$
It now follows that $u=\mathbf{R}[f,g]$ has the correct boundary data along the axes $x=0$ and $y=0$.
Theorem 1.2.
(Riemann) Provided that $f,g$ are continuous univariate functions with $f(0)=g(0)$, the function $u(x,y)=\mathbf{R}[f,g](x,y)$ is a continuous function of two variables, which solves the Darboux-Goursat problem (1.2) in the sense of distribution theory.
To offer a cohesive reader-friendly presentation, we supply the proof of this theorem in the appendix (Section 7). Unfortunately, we failed to pinpoint the exact reference to Riemann’s contribution here in his collected works [23] (perhaps the paper Ueber die Fortplantzung ebener Luftwellen von endlicher Schwingungsweite is the most relevant), but it is generally accepted that the method, which is mainly concerned with the Darboux-Goursat problem for inhomogeneous equations, is due to Riemann, and thus called the Riemann method.
It remains to determine that Riemann’s solution is the unique solution to the Darboux-Goursat problem (1.2). To this end, it is an important observation that continuous distributional solutions to the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ are characterized by the quadrature identity
(1.6) $\int_{a}^{a^{\prime}}\int_{b}^{b^{\prime}}u(s,t)\,\mathrm{d}t\mathrm{d}s=u(a^{\prime},b)+u(a,b^{\prime})-u(a^{\prime},b^{\prime})-u(a,b).$
If we apply (1.6) to $a^{\prime}=x,b^{\prime}=y,a=b=0$, we may view the equation as a way to represent $u(x,y)$ in terms of $u(x,0)$, $u(0,y)$, $u(0,0)$, and the integral over the corresponding rectangle:
(1.7) $u(x,y)=u(x,0)+u(0,y)-u(0,0)-\int_{0}^{x}\int_{0}^{y}u(s,t)\,\mathrm{d}t\mathrm{d}s.$
Based on this quadrature identity, the Picard method applies to give uniqueness.
Theorem 1.3.
(Picard’s method) Suppose $u(x,y)$ is a continuous solution of the Darboux-Goursat problem (1.2), where the Klein-Gordon equation is understood in the sense of distribution theory, and the univariate functions $f,g$ are continuous with $f(0)=g(0)$. Then $u$ coincides with Riemann’s solution: $u=\mathbf{R}[f,g]$.
The proof of this known theorem is supplied for completeness in the appendix (Section 7).
1.2. Finite speed propagation and uniqueness
The Klein-Gordon equation is relativistic in the sense that the speed of light $c$ is a finite positive constant, which in our modelling is set to $c=1$. Let us see what this means in terms of Riemann’s solution $u=\mathbf{R}[f,g]$. So, we suppose that the continuous data $f,g$ in the Darboux-Goursat problem (1.2) vanishes initially:
$f(t)=0\,\,\,\text{for all}\,\,\,t\in[a,a^{\prime}],\quad g(t)=0\,\,\,\text{for all}\,\,\,t\in[b,b^{\prime}],$
for some reals $a\leq 0\leq a^{\prime}$ and $b\leq 0\leq b^{\prime}$. It then follows by inspection of Riemann’s solution that
$u(x,y)=\mathbf{R}[f,g](x,y)=0,\qquad(x,y)\in[a,a^{\prime}]\times[b,b^{\prime}].$
This has the interpretation that information travels at finite speed, and that the values on the rectangle $a,a^{\prime}]\times[b,b^{\prime}]$ are fully determined by the values on the union of the characteristic segments $a,a^{\prime}]\times\{0\}$ and $\{0\}\times[b,b^{\prime}]$. Nevertheless, this presupposes that $u$ is a global continuous solution to the Klein-Gordon equation. What if $u$ is just a continuous function on the rectangle $a,a^{\prime}]\times[b,b^{\prime}]$ that that solves the Klein-Gordon equation in the interior in the sense of distribution theory? The Picard argument still applies to give uniqueness:
Theorem 1.4.
(Picard’s method) Suppose $a\leq 0\leq a^{\prime}$ and $b\leq 0\leq b^{\prime}$, and that $u$ is a continuous function on the closed rectangle $a,a^{\prime}]\times[b,b^{\prime}]$ which solves the Klein-Gordon equation in the interior of the rectangle in the sense of distribution theory. If
$u(x,0)=u(0,y)=0,\qquad x\in[a,a^{\prime}],\,\,\,y\in[b,b^{\prime}],$
then $u=0$ holds on $a,a^{\prime}]\times[b,b^{\prime}]$.
The proof of this theorem is basically the same as that of Theorem 1.3. It can be found in the appendix (Section 7).
1.3. Poincaré group invariance of the Klein-Gordon equation
We formalize the invariance of the Klein-Gordon equation under the Poincaré group in a proposition.
Proposition 1.5.
For a real parameter $\lambda\neq 0$, and two real parameters $a,b$, let $L$ and $M$ denote the affine transformations $L(x,y)=(\lambda x+a,\lambda^{-1}y+b)$ and $M(x,y)=(\lambda y+a,\lambda^{-1}x+b)$. If $u$ solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ on a domain $\Omega\subset\mathbb{R}^{2}$, then so do $v=u\circ L$ and $w=u\circ M$ on the corresponding domains $L^{-1}(\Omega)$ and $M^{-1}(\Omega)$, respectively.
This result is well-known and we omit the proof.
Remark 1.6.
The assertion of the proposition holds not only for classical solutions but for (weak) solutions in the sense of distribution theory as well.
1.4. Gluing solutions of the Klein-Gordon equation along characteristics
Suppose we are given a bounded convex domain $\Omega\in\mathbb{R}^{2}$, and a line $L\subset\mathbb{R}^{2}$ which is such that $\Omega\cap L$ has positive length. Suppose moreover $u:\Omega\to\mathbb{C}$ is continuous and harmonic in $\Omega\setminus L$. Then it is not necessarily true that $u$ extends to be harmonic in $\Omega$. What can prevent harmonic extension is the appearance of a Neumann jump along $L$, that is, the first-order derivative along the the normal direction has a jump across $L$. As we replace the Laplacian (which has order $2$) with Cauchy-Riemann operator $\bar{\partial}$, which has order $1$, this problem does not occur, and we automatically get analytic continuation along any given linear segment $L\cap\Omega$ if the given function is continuous on $\Omega$ and holomorphic in $\Omega\setminus L$. As the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ has order $2$, the appearance of a nontrivial Neumann jump along the line $L$ may influence the extendability across the line as a solution. This, however, is not the case if the line is characteristic.
Theorem 1.7.
Suppose $\Omega\subset\mathbb{R}^{2}$ is bounded and convex and $L$ is a line which is characteristic, that is, parallel to either the $x$-axis or the $y$-axis. If $u:\Omega\to\mathbb{C}$ is continuous and solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in $\Omega\setminus L$ in the sense of distribution theory, then $u$ extends to a solution to the Klein-Gordon equation on all of $\Omega$.
The proof of this (likely known) theorem is supplied in the appendix (Section 7).
1.5. Splitting Riemann’s solution into unilateral wave solutions
Riemann’s solution $\mathbf{R}[f,g]$ to the Darboux-Goursat problem (1.2) may be split into three components.
Proposition 1.8.
If $u=\mathbf{R}[f,g]$ is Riemann’s solution to the Darboux-Goursat problem (1.2) with continuous data $f,g$, subject to the compatibility condition $f(0)=g(0)$, then $u$ splits as
$u=u_{0}+u_{\mathrm{horz}}+u_{\mathrm{vert}},$
where
$u_{0}(x,y)=f(0)J_{0,0}(x,y)=f(0)\mathbf{R}[1,1](x,y)$
and
$u_{\mathrm{horz}}=\mathbf{R}[f_{0},0],\quad u_{\mathrm{vert}}=\mathbf{R}[0,g_{0}],$
with $f_{0}(t)=f(t)-f(0)$ and $g_{0}(t)=g(t)-g(0)$. Here, each component $u_{0}$, $u_{\mathrm{horz}}$, $u_{\mathrm{vert}}$ is continuous and solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ individually.
This assertion is immediate from Theorems 1.2 and by slightly rearranging the terms in Riemann’s solution. Alternatively, we can observe that the expressed solution $u$ solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory since each of the terms $u_{0},u_{\mathrm{horz}},u_{\mathrm{vert}}$ does, and that the values along the the $x$-axis and the $y$-axis coincide with $f$ and $g$, respectively. Theorem 1.3 then shows that $u=u_{0}+u_{\mathrm{horz}}+u_{\mathrm{vert}}$ must coincide with $\mathbf{R}[f,g]$. Proposition 1.8 motivates the following terminology.
Definition 1.9.
The wave $u_{\mathrm{horz}}=\mathbf{R}[f_{0},0]$ is called a horizontal unilateral wave solution, whereas the wave $u_{\mathrm{vert}}=\mathbf{R}[0,g_{0}]$ is called a vertical unilateral wave solution. Together, we refer to the waves of the form $u_{\mathrm{horz}}$ or $u_{\mathrm{vert}}$ as unilateral wave solutions.
The following proposition characterizes such unilateral wave solutions.
Proposition 1.10.
Suppose $u:\mathbb{R}^{2}\to\mathbb{C}$ is continuous and solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory. Then
(a) $u=u_{\mathrm{horz}}$ holds if and only if $u(0,y)=0$ for all $y\in\mathbb{R}$, and
(b) $u=u_{\mathrm{vert}}$ holds if and only if $u(x,0)=0$ for all $x\in\mathbb{R}$.
In view of the properties of Riemann’s solution $u=\mathbf{R}[f,g]$ outlined in Theorems 1.2 and 1.3, the assertion of the proposition is immediate from the definition of unilateral wave solutions, horizontal and vertical.
It is the aim of this paper to study the growth properties of the unilateral wave solutions $u_{\mathrm{horz}}$ and $u_{\mathrm{vert}}$. By Lorentz group invariance, it is equivalent to focus on just one of them, and we will consider the horizontal unilateral wave solution $u_{\mathrm{horz}}$. As it turns out, there is Liouville phenomenon for such solutions. Here, we recall the classical theorem of Liouville, which asserts that any entire function which is also bounded has to be constant. There are many extensions, for instance, a harmonic function in the entire plane which is bounded must be constant as well. A version which is particularly attractive is known as the Phragmén-Lindelöf principle which asserts that if a harmonic function vanishes on a line and has insufficient growth on one side of the line, then it must vanish everywhere.
1.6. Motivation: the Phragmén-Lindelöf principle for harmonic functions
The analogous instance of the Laplace equation in place of the Klein-Gordon equation is well-known, and called the Phragmén-Lindelöf principle.
Theorem 1.11.
[…]
$u(x,y)=\mathrm{o}(y),\qquad(x,y)\in\bigcup_{n=1}\partial D_{n},$
holds as $|(x,y)|\to+\infty$. Then $u(x,y)\equiv 0$ identically.
Remark 1.12.
The proof of the the Phragmén-Lindelöf principle is based on the maximum principle, which applies to the real and imaginary parts of $u$ separately. The example $u(x,y)=y$ shows that the growth condition on the harmonic function $u$ optimal. As for the Klein-Gordon equation, we cannot access any maximum principle directly, given that the equation is hyperbolic.
The connection with Liouville’s theorem is actually rather direct: By Schwarz reflection, which applies along the linear boundary since the values of the harmonic function vanish there, the harmonic function $u$ extends to a harmonic function in the whole complex plane with $u(\bar{z})=-u(z)$. The growth assumptions on this entire harmonic function are such that we have very slow growth. In other words, we are close in spirit to the classical Liouville theorem.
Remark 1.13.
There is a rather sizeable literature on Liouville type phenomena for elliptic equations, linear and nonlinear, also involving fractional Laplacians, etc. We could mention a particularly interesting effect in the discrete context, where harmonicity combined with boundedness on a big enough portion is enough to force the function to be constant [8]. This is in contrast with the continuous setting where it is enough to have a narrow path tending to infinity to permit the existence of a nontrivial harmonic function which grows wildly along the path while it is bounded elsewhere.
1.7. Liouville type phenomenon for spacelike quarter-planes
Apparently, in the context of the Phragmén-Lindelöf principle, there is some degree of minimal growth which must be surpassed to get a nontrivial solution. What about the unilateral solutions to the Klein-Gordon equation? By the characterization of Proposition 1.10, they vanish either on the horizontal line $\mathbb{R}\times\{0\}$ or the vertical line $\{0\}\times\mathbb{R}$. This may be able to force them to grow somewhere, and, in fact, as it turns out, the required growth is rather strong. A first result of this kind which we may present is the following. Note that with respect to the origin $(0,0)$, the equation $xy=0$ represents the light cone emanating from $(0,0)$, while $xy>0$ represents the timelike vectors, and $xy<0$ represents the spacelike vectors. The spacelike vectors can all be thought of as representing events contemporary with that of the origin.
Theorem 1.14.
Suppose $u$ is a continuous complex-valued function on the closed spacelike quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, which solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory in the interior quarter-plane $\mathbb{R}_{>0}\times\mathbb{R}_{<0}$. Suppose moreover that $u$ vanishes along the semi-axis $\{0\}\times\mathbb{R}_{\leq 0}$. Finally, we assume that $u$ meets the uniform growth bound
$|u(x,y)|=\mathrm{O}\big(\exp(\theta_{1}x+\theta_{2}|y|)\big),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathcal{Y},$
where $\theta_{1},\theta_{2}\in\mathbb{R}_{>0}$ and the nonempty subset $\mathcal{Y}\subset\mathbb{R}_{\leq 0}$ has $0\in\mathcal{Y}$ and $\inf\mathcal{Y}=$-$\infty$. Then, if $\theta_{1}\theta_{2}<1$, it follows that the solution collapses to $u(x,y)=0$ for all $(x,y)\in\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$.
The proof of Theorem 1.14 is supplied in Subsection 2.3.
Remark 1.15.
(a) If we compare with Theorem 1.11, the union of lines $\mathbb{R}_{\geq 0}\times\mathcal{Y}$ corresponds to the union of fronts $\cup_{n}\partial D_{n}$.
(b) The fact that the criterion involves the product $\theta_{1}\theta_{2}$ is a consequence of the Lorentz group invariance of the Klein-Gordon equation (see Proposition 1.5).
(c) In the instance that $\mathcal{Y}=\mathbb{R}_{\leq 0}$, it appears that Theorem 1.14 was first obtained by Andrew Bakan (unpublished). After all, he mentioned a version of it in a seminar presentation in Stockholm during his visit in the fall semester of 2016. We have since learned that Bakan’s interest in the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ has its roots in the 1990s.
The condition that $u$ vanishes along the semi-axis $\mathbb{R}_{\geq 0}\times\{0\}$ entails that $u$ may be extended to a solution to the Klein-Gordon equation in the whole plane $\mathbb{R}^{2}$ which vanishes on the line $\{0\}\times\mathbb{R}$, that is, the extension is a horizontal unilateral solution.
It is natural to wonder about the sharpness of the conditions in Theorem 1.14. To this end, we have the following.
Theorem 1.16.
Suppose $\theta_{1},\theta_{2}\in\mathbb{R}_{>0}$ with $\theta_{1}\theta_{2}\geq 1$. There exists a nontrivial continuous horizontal unilateral solution $u$ to the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ such that the restriction of $u$ to the quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$ does not vanish identically, while it meets the uniform growth condition
$|u(x,y)|=\mathrm{O}\big(\exp(\theta_{1}x+\theta_{2}|y|)\big),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}.$
The proof of Theorem 1.16 is supplied in Subsection 2.1.
It is a natural question to ask if we can play with the growth conditions while keeping the conclusion of Theorem 1.14. For instance, if the growth in the $x$ direction is slower, can we allow for faster growth in the $y$ direction? It turns out that this question has answers which connect with the foundations of harmonic analysis, and more precisely, with the issue of so-called analytic non-quasianalyticity, as studied by Arne Beurling [5].
Theorem 1.17.
Suppose $u$ is a continuous complex-valued function on the closed spacelike quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, which solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory in the interior quarter-plane $\mathbb{R}_{>0}\times\mathbb{R}_{<0}$. Suppose moreover that $u$ vanishes along the semi-axis $\{0\}\times\mathbb{R}_{\leq 0}$. Finally, we assume that the nonempty subset $\mathcal{Y}\subset\mathbb{R}_{\leq 0}$ has $\inf\mathcal{Y}=$-$\infty$, and that $u$ enjoys the growth bound
$\forall y\in\mathcal{Y}:\,\,\,|u(x,y)|=\mathrm{O}_{y}\big(\exp(\theta x^{q})\big),$
as $x\to+\infty$, where the implied constant is allowed to depend on $y$, and the parameters $q,\theta$ are confined to $0<q<\frac{1}{2}$ and $0<\theta<+\infty$. It then follows that the solution collapses to $u=0$ on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$.
The proof of Theorem 1.17 is based on the work of Hayman and Korenblum [14], [15], and on Korenblum’s ICM address [22], which in turn mentions the connection with Beurling’s contribution. It is supplied in Subsection 4.3.
In Theorem 1.17, we do not require any particular growth control on $u(x,y)$ in the negative $y$ direction, and the price we pay is a rather strict growth bound in the positive $x$ direction. In Theorem 1.14, on the other hand, we have basically equal control in both directions. There is a compromise which produces intermediate results. To formulate it, we need the following concept, which requires a subset along the semi-axis to be evenly spread-out according to a certain density.
Definition 1.18.
For $\frac{1}{2}<q<1$ and $q^{\prime\prime}:=q/(2q-1)$, a subset $\mathcal{Y}\subset\mathbb{R}_{\leq 0}$ is said to induce an asymptotic $q$-covering of $\mathbb{R}_{\leq 0}$ if there exists a constant $M\in\mathbb{R}_{>0}$ such that
$\infty,y_{1}]\subset\bigcup_{y\in\mathcal{Y}}[y-MR(y,q),y+MR(y,q)]$
holds for some $y_{1}\in\mathbb{R}_{<0}$, where
$R(y,q):=|y|\,\min\{1,|y|^{-\frac{q^{\prime\prime}}{2}}\}.$
Remark 1.19.
For $\frac{1}{2}<q<1$ we know that $1<q^{\prime\prime}<+\infty$. For $q$ close to $1$, $q^{\prime\prime}$ gets close to $1$, so that this condition allows for rather sparse sets $\mathcal{Y}$, whereas for $q$ close to $\frac{1}{2}$, $q^{\prime\prime}$ gets close to $\infty$, and then the set $\mathcal{Y}$ has to be much denser.
Theorem 1.20.
Suppose $u$ is a continuous complex-valued function in the quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, which solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory in the interior quarter-plane $\mathbb{R}_{>0}\times\mathbb{R}_{<0}$. Suppose moreover that $u$ vanishes along the semi-axis $\{0\}\times\mathbb{R}_{\leq 0}$. Finally, we assume that $u$ meets the uniform growth bound
$|u(x,y)|=\mathrm{O}\big(\exp(\theta_{1}|x|^{q}+\theta_{2}|y|^{q^{\prime\prime}})\big),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathcal{Y},$
where $0<\theta_{1},\theta_{2}<+\infty$, $\frac{1}{2}<q<1$ and $q^{\prime\prime}=q/(2q-1)$, and $0\in\mathcal{Y}$, where the subset $\mathcal{Y}\subset\mathbb{R}_{\leq 0}$ is supposed to induce an asymptotic $q$-covering of $\mathbb{R}_{\leq 0}$. If we have the bound
$\theta_{1}^{1/q}\theta_{2}^{1/q^{\prime\prime}}<\bigg(\frac{1}{q}\bigg)^{1/q}\bigg(\frac{1}{q^{\prime\prime}}\bigg)^{1/q^{\prime\prime}}\cos^{2}\frac{\pi(1-q)}{2q},$
it then follows that the solution collapses to $u=0$ on all of $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$.
The proof of this theorem is supplied in Section 5.
Remark 1.21.
We believe that the result is essentially sharp. In particular, if the reverse inequality
$\theta_{1}^{1/q}\theta_{2}^{1/q^{\prime\prime}}>\bigg(\frac{1}{q}\bigg)^{1/q}\bigg(\frac{1}{q^{\prime\prime}}\bigg)^{1/q^{\prime\prime}}\cos^{2}\frac{\pi(1-q)}{2q},$
holds, then there should exist counterexamples. Here, Borichev’s work [6] could be a starting point, which in fact elaborates on an earlier pioneering insight by Mary Cartwright [10].
In terms of the parameter $q$ appearing in Theorems 1.17 and 1.20, the range $0<q<\frac{1}{2}$ is covered by Theorem 1.17, while the range $\frac{1}{2}<q<1$ constitutes Theorem 1.20, and $q=1$ amounts to Theorem 1.14. The value $q=\frac{1}{2}$ is not covered by any of these results, so it needs to be treated separately.
Theorem 1.22.
Suppose $u$ is a continuous complex-valued function in the quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, which solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory in the interior quarter-plane $\mathbb{R}_{>0}\times\mathbb{R}_{<0}$. Suppose moreover that $u$ vanishes along the semi-axis $\{0\}\times\mathbb{R}_{\leq 0}$. Finally, we assume that $u$ meets the uniform growth bound
$|u(x,y)|=\mathrm{O}\big(\exp\big(\theta_{1}\sqrt{|x|}+\exp(\theta_{2}\sqrt{|y|})\big)\big),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0},$
where the parameters $\theta_{1},\theta_{2}\in\mathbb{R}_{>0}$ satisfy $\theta_{1}\theta_{2}<2\pi$. It then follows that the solution collapses to $u=0$ on all of $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$.
The proof of this theorem is supplied in Section 6.
Remark 1.23.
(a) It is possible to replace the whole half-line $\mathbb{R}_{\leq 0}$ by a discrete subset $\mathcal{Y}$, subject to a density condition, along the lines of Theorem 1.20. This time, the distance between consecutive points in $\mathcal{Y}$ should be on the order of magnitude $\exp(-\frac{1}{2}\theta_{2}|y|^{-\frac{1}{2}})$ for large negative $y$.
(b) We expect the parameter range $\theta_{1}\theta_{2}<2\pi$ to be sharp, in the sense that for $\theta_{1}\theta_{2}>2\pi$, it should be possible to find counterexamples.
Remark 1.24.
The remaining parameter range $1<q<+\infty$ is beyond the treatment with the standard Laplace transform, which is our main tool here. However, it may be possible to use in its place an approximate Laplace transform like in the theory of almost holomorphic functions due to Evseĭ Dyn’kin, and used later by, e.g., Volberg [24], and Borichev and Hedenmalm [7].
1.8. Acknowledgements
As mentioned earlier, this work continues early insights by Andrew Bakan, dating back many years. We definitely appreciate his contributions to the topic.
2. Horizontal unilateral wave solutions
2.1. An example of growth in a spacelike quadrant
We consider the Riemann solution $u_{1}=\mathbf{R}[f_{1},0]$, where the continuous function $f_{1}$ is given by
$f_{1}(x):=\begin{cases}0,\qquad x\leq 0,\\ x,\qquad 0<x<+\infty.\\ 1,\qquad x\geq 1.\end{cases}$
Then $u_{1}$ is a horizontal unilateral wave solution, expressed by
$u_{1}(x,y)=\mathbf{R}[f_{1},0](x,y)=f_{1}(x)-\int_{0}^{x}J_{0,1}(x-s,y)f_{1}(s)\,\mathrm{d}s.$
Since we study $y\in\mathbb{R}_{\leq 0}$, it is convenient to flip $y\mapsto-y$:
(2.1) $u_{1}(x,-y)=\mathbf{R}[f_{1},0](x,-y)=f_{1}(x)-\int_{0}^{x}J_{0,1}(x-s,-y)f_{1}(s)\,\mathrm{d}s\\ =f_{1}(x)-\int_{0}^{x}\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!(j+1)!}(x-s)^{j}(-y)^{j+1}f_{1}(s)\,\mathrm{d}s\\ =f_{1}(x)+\sum_{j=0}^{+\infty}\frac{y^{j+1}}{j!(j+1)!}\int_{0}^{x}(x-s)^{j}f_{1}(s)\,\mathrm{d}s.$
Since $0\leq f_{1}(x)\leq 1$ holds for $x\in\mathbb{R}_{\geq 0}$, it now follows that
(2.2) $0\leq u_{1}(x,-y)\leq 1+\sum_{j=0}^{+\infty}\frac{y^{j+1}}{j!(j+1)!}\int_{0}^{x}(x-s)^{j}\mathrm{d}s\\ =1+\sum_{j=0}^{+\infty}\frac{(xy)^{j+1}}{[(j+1)!]^{2}}=J_{0,0}(x,-y),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}.$
Next, the function $J_{0,0}$ may be expressed in terms of the standard Bessel function:
$J_{0,0}(x,-y)=J_{0}(2\sqrt{-xy}),$
Classical asymptotics of the Bessel function gives that as $xy\to+\infty$ while (x,y)(x,-y) is in the spacelike quarter-plane where $x,y>0$, we have that
(2.3) $J_{0,0}(x,-y)=(1+\mathrm{o}(1))\frac{1}{2\sqrt{\pi}}\,(xy)^{-1/4}\exp\big(2\sqrt{xy}\big),$
whereas if $xy$ is kept bounded, we have
$J_{0,0}(x,-y)=\mathrm{O}(1).$
Proof of Theorem 1.16.
By Young’s inequality for products, in fact, in this case, just by completing squares, we have that
$2\sqrt{xy}\leq\theta_{1}x+\theta_{1}^{-1}y\leq\theta_{1}x+\theta_{2}y,\qquad x,y\in\mathbb{R}_{\geq 0},$
given that $\theta_{1}\theta_{2}\geq 1$ entails that $\theta_{2}\geq\theta_{1}^{-1}$. It now follows from the above control on $J_{0,0}(x,-y)$ combined with the inequalities in (2.2) that
$u_{1}(x,-y)=\mathrm{O}(\exp(\theta_{1}x+\theta_{2}y)),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0},$
holds uniformly. Moreover, since $u_{1}(x,0)=f_{1}(x)$ holds for $x\in\mathbb{R}$, it follows that $u_{1}$ is the nontrivial solution to the Klein-Gordon equation which is in particular does not vanish identically on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$ and enjoys the claimed growth estimate. ∎
2.2. The evolution equation for horizontal unilateral wave solutions
We consider a general horizontal unilateral wave solution $u_{\mathrm{horz}}=\mathbf{R}[f,0]$, where $f$ is a continuous function on the line with $f(0)=0$. Explicitly, $u_{\mathrm{horz}}$ is given by
(2.4) $u_{\mathrm{horz}}(x,y)=f(x)-\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s.$
The formula (2.4) is of convolution type on the half-line, which suggests that we might try to analyze it using the classical Laplace transform. To this end, we assume that $f$ does not grow faster than exponentially in the mean. More explicitly, we require that
$\forall\sigma>\sigma_{0}:\quad\int_{0}^{+\infty}\mathrm{e}^{-x\sigma}|f(x)|\mathrm{d}x<+\infty,$
holds for some parameter value $\sigma_{0}\in\mathbb{R}_{\geq 0}$. Then the Laplace transform $\mathbf{L}[f]$ given by
$\mathbf{L}[f](\zeta)=\int_{0}^{+\infty}\mathrm{e}^{-x\zeta}f(x)\,\mathrm{d}x,\qquad\mathrm{Re}\,\zeta>\sigma_{0},$
defines a holomorphic function in the given half-plane, and if we write
$J_{0,1}(\cdot,y)\ast f](x)=\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s$
for the semigroup convolution, the Laplace transform of the convolution is just the product of the two individual Laplace transforms,
$\mathbf{L}[J_{0,1}(\cdot,y)\ast f](\zeta)=\mathbf{L}[J_{0,1}(\cdot,y)](\zeta)\mathbf{L}[f](\zeta),$
where
$\mathbf{L}[J_{0,1}(\cdot,y)](\zeta)=\int_{0}^{+\infty}\mathrm{e}^{-x\zeta}J_{0,1}(x,y)\,\mathrm{d}x=1-\mathrm{e}^{-y/\zeta},\qquad\mathrm{Re}\,\zeta>0.$
Although known, this calculation is surprisingly hard to find in the literature. It is, however, rather straightforward:
(2.5) $\int_{0}^{+\infty}\mathrm{e}^{-x\zeta}J_{0,1}(x,y)\,\mathrm{d}x=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!(j+1)!}\int_{0}^{+\infty}x^{j}\,y^{j+1}\mathrm{e}^{-x\zeta}\mathrm{d}x\\ =\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{(j+1)!}y^{j+1}\zeta^{-j-1}=1-\mathrm{e}^{-y/\zeta},$
again provided that $\mathrm{Re}\,\zeta>0$. It is now immediate from (2.4) that
(2.6) $\mathbf{L}[u_{\mathrm{horz}}(\cdot,y)](\zeta)=\int_{0}^{+\infty}\mathrm{e}^{-x\zeta}u_{\mathrm{horz}}(x,y)\,\mathrm{d}x=\mathbf{L}[f](\zeta)-\mathbf{L}[J_{1,0}(y,\cdot)\ast f](\zeta)\\ =\mathbf{L}[f](\zeta)-\mathbf{L}[J_{1,0}(y,\cdot)](\zeta)\mathbf{L}[f](\zeta)=\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta),\qquad\mathrm{Re}\,\zeta>\sigma_{0}.$
The equation (2.6) is fundamental to our analysis of the solutions $u_{\mathrm{horz}}=\mathbf{R}[f,0]$. The evolution in the $y$ parameter is simple, and indeed, it resembles the pattern used in the scattering method to analyze certain nonlinear integrable partial differential equations such as the KdV (Korteweg-de Vries) equation. Basically, we perform the so-called scattering transformation, and afterwards apply a smooth flow evolution, and then get back to the original setting via inverse scattering. For the basics on the method, see, e.g., the books [4], [1].
2.3. Uniqueness in the exponential growth case $q=1$
We turn to the proof of the simplest instance of the Liouville-type phenomenon for the Klein-Gordon equation.
Proof of Theorem 1.14.
We let $f(x):=u(x,0)$ for $\in\mathbb{R}_{\geq 0}$, and observe that by continuity, we must have $f(0)=0$. Next, Riemann’s solution $\mathbf{R}[f,0]$ is well-defined in the half-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}$, and vanishes on the half-line $\{0\}\times\mathbb{R}$. Then the difference function $u-\mathbf{R}[f,0]$ is continuous, solves the Klein-Gordon equation in the sense of distribution theory, and vanishes along the two characteristic semi-axes $\mathbb{R}_{\geq 0}\times\{0\}$ and $\{0\}\times\mathbb{R}_{\leq 0}$. By Theorem 1.4, then, $u-\mathbf{R}[f,0]=0$ on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, so that $u=\mathbf{R}[f,0]$ holds there, so that we may think of $u=u_{\mathrm{horz}}$ as a unilateral horizontal wave solution. The given growth control on $u=\mathbf{R}[f,0]$ says that
$|u(x,y)|\leq C_{0}\exp(\theta_{1}x+\theta_{2}|y|),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathcal{Y},$
for some positive constant $C_{0}$. In particular, since $0\in\mathcal{Y}$, we have
$|f(x)|=|u(x,0)|\leq C_{0}\exp(\theta_{1}x),\qquad x\in\mathbb{R}_{\geq 0}.$
The Laplace transform
$\mathbf{L}[f](\zeta)=\int_{0}^{+\infty}\mathrm{e}^{-t\zeta}f(t)\,\mathrm{d}t,\qquad\mathrm{Re}\,\zeta>\theta_{1},$
is then holomorphic in the indicated half-plane, with the growth bound
$|\mathbf{L}[f](\zeta)|\leq\frac{C_{0}}{\mathrm{Re}\,\zeta-\theta_{1}},\qquad\mathrm{Re}\,\zeta>\theta_{1}.$
Given the growth control on the solution $u=u_{\mathrm{horz}}=\mathbf{R}[f,0]$, we find analogously that
$|\mathbf{L}[u(\cdot,y)](\zeta)|\leq\frac{C_{0}\exp(\theta_{2}|y|)}{\mathrm{Re}\,\zeta-\theta_{1}},\qquad\mathrm{Re}\,\zeta>\theta_{1},\,\,\,y\in\mathcal{Y}.$
Finally, in view of the evolution equation for the Laplace transform (2.6), we know that
$|\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta)|\leq\frac{C_{0}\exp(\theta_{2}|y|)}{\mathrm{Re}\,\zeta-\theta_{1}},\qquad\mathrm{Re}\,\zeta>\theta_{1},\,\,\,y\in\mathcal{Y},$
and hence
$|\mathbf{L}[f](\zeta)|\leq\frac{C_{0}}{\mathrm{Re}\,\zeta-\theta_{1}}\,\exp\big(\theta_{2}|y|+\mathrm{Re}\,(y/\zeta)\big),\qquad\mathrm{Re}\,\zeta>\theta_{1}.$
The left-hand side does not depend on $y\in\mathcal{Y}$, so we are free to minimize over $y$. Since y=|y|y=-|y|, we find that
$|\mathbf{L}[f](\zeta)|\leq\frac{C_{0}}{\mathrm{Re}\,\zeta-\theta_{1}}\,\inf_{y\in\mathcal{Y}}\,\exp\big(|y|(\theta_{2}-\mathrm{Re}\,(1/\zeta))\big),\qquad\mathrm{Re}\,\zeta>\theta_{1},\,\,\,y\in\mathcal{Y},$
which gives that
$|\mathbf{L}[f](\zeta)|=0\quad\text{if}\,\,\,\mathrm{Re}\,\zeta>\theta_{1}\,\,\,\text{and}\,\,\,\mathrm{Re}\,(1/\zeta)>\theta_{2}.$
Given that $\theta_{1}\theta_{2}<1$, the indicated region of vanishing is open and nonempty, so that by the uniqueness theorem for holomorphic functions, it follows that $\mathbf{L}[f]=0$ identically, an hence $f=0$ on $\mathbb{R}_{\geq 0}$. Consequently, the solution $u=\mathbf{R}[f,0]$ vanishes in the half-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}$, and, in particular, in the quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$. ∎
3. An auxiliary integral estimate
3.1. Estimation of an integral
The following estimate is probably known but we have found no suitable reference.
Lemma 3.1.
Suppose that $0<q<1$ and $0<\theta,\sigma<+\infty$. Then
(3.1) $\int_{0}^{+\infty}\exp(\theta\,t^{q}-\sigma t)\,\mathrm{d}t\leq\frac{1}{\sigma}\big(A(q,\theta)\sigma^{-\frac{q}{2(1-q)}}+B(q)\big)\,\exp\big(D(q,\theta)\,\sigma^{-\frac{q}{1-q}}\big),$
where
$A(q,\theta)=2^{1-\frac{q}{2}}(q\theta)^{\frac{1}{2(1-q)}}\sqrt{\frac{2\pi}{1-q}}\,,\quad B(q)=\frac{4}{1-q},\quad D(q,\theta)=(1-q)\theta(q\theta)^{\frac{q}{1-q}}.$
The key estimate to arrive at Lemma 3.1 is the following.
Lemma 3.2.
Suppose $\psi:\mathbb{R}_{\geq 0}\to\mathbb{R}$ is $C^{3}$-smooth on $\mathbb{R}_{>0}$, with $\psi^{\prime\prime}>0$ and $\psi^{\prime\prime\prime}<0$ on $\mathbb{R}_{>0}$. If $t_{0}\in\mathbb{R}_{>0}$ is such that $\psi^{\prime}(t_{0})=0$, and $0<t_{0}<t_{1}<+\infty$, we then have the estimate
(3.2) $\int_{0}^{+\infty}\exp(-\psi(t))\,\mathrm{d}t\leq\mathrm{e}^{-\psi(t_{0})}\sqrt{\frac{2\pi}{\psi^{\prime\prime}(t_{1})}}+\frac{\mathrm{e}^{-\psi(t_{1})}}{\psi^{\prime}(t_{1})}.$
Proof.
We begin by splitting the integral:
(3.3) $\int_{0}^{+\infty}\exp(-\psi(t))\,\mathrm{d}t=\int_{0}^{t_{1}}\exp(-\psi(t))\,\mathrm{d}t+\int_{t_{1}}^{+\infty}\exp(-\psi(t))\,\mathrm{d}t.$
Since $\psi^{\prime\prime}>0$, the function $\psi$ is strictly convex, and consequently,
(3.4) $\psi(t)\geq\psi(t_{1})+(t-t_{1})\psi^{\prime}(t_{1}),\qquad x\in\mathbb{R}_{\geq 0},$
since convex functions have support lines at each point. Moreover, since $\psi^{\prime}(t_{0})=0$ and $0<t_{0}<t_{1}<+\infty$, and $\psi^{\prime\prime}>0$, we know that $\psi^{\prime}(t_{1})>0$. As regards the second integral on the right-hand side of (3.3), it now follows that
(3.5) $\int_{t_{1}}^{+\infty}\exp(-\psi(t))\,\mathrm{d}t\leq\int_{t_{1}}^{+\infty}\exp\big(-\psi(t_{1})-(t-t_{1})\psi^{\prime}(t_{1})\big)\mathrm{d}t=\frac{\mathrm{e}^{-\psi(t_{1})}}{\psi^{\prime}(t_{1})}.$
Regarding the first integral on the right-hand side of (3.3), we use the Lagrange form of the remainder in Taylor’s formula:
(3.6) $\psi(t)=\psi(t_{0})+(t-t_{0})\psi^{\prime}(t_{0})+\frac{1}{2}(t-t_{0})^{2}\psi^{\prime\prime}(\xi)=\psi(t_{0})+\frac{1}{2}(t-t_{0})^{2}\psi^{\prime\prime}(\xi),$
where $\xi$ is an intermediate point between $t_{0}$ and $t$. Since $\psi^{\prime\prime\prime}<0$, the second derivative $\psi^{\prime\prime}$ is strictly decreasing, so that
(3.7) $\psi(t)=\psi(t_{0})+\frac{1}{2}(t-t_{0})^{2}\psi^{\prime\prime}(\xi)\geq\psi(t_{0})+\frac{1}{2}(t-t_{0})^{2}\psi^{\prime\prime}(t_{1}),\qquad 0<t\leq t_{1}.$
It now follows that
(3.8) $\int_{0}^{t_{1}}\exp(-\psi(t))\,\mathrm{d}t\leq\int_{0}^{t_{1}}\exp\bigg(-\psi(t_{0})-\frac{1}{2}(t-t_{0})^{2}\psi^{\prime\prime}(t_{1})\bigg)\mathrm{d}t\\ \leq\mathrm{e}^{-\psi(t_{0})}\int_{-\infty}^{+\infty}\exp\bigg(-\frac{1}{2}(t-t_{0})^{2}\psi^{\prime\prime}(t_{1})\bigg)\mathrm{d}t=\mathrm{e}^{-\psi(t_{0})}\sqrt{\frac{2\pi}{\psi^{\prime\prime}(t_{0})}}.$
The claimed estimate is a consequence of the decomposition (3.3) and the estimates (3.5) and (3.8), which completes the proof. ∎
Proof of Lemma 3.1.
If $\psi(t)=\sigma t-\theta t^{q}$, we have $\psi^{\prime\prime}(t)=\theta q(1-q)t^{q-2}>0$ and $\psi^{\prime\prime\prime}(t)=-\theta q(1-q)(2-q)t^{q-3}<0$, and the point $t_{0}$ with $\psi^{\prime}(t_{0})=0$ is given by
$t_{0}=\bigg(\frac{\theta q}{\sigma}\bigg)^{\frac{1}{1-q}},$
so that
$\psi(t_{0})=-(1-q)\theta(q\theta/\sigma)^{\frac{q}{1-q}}.$
If $t_{1}>t_{0}$, then there is a point $\xi$ between $t_{0}$ and $t_{1}$ such that
$\psi^{\prime}(t_{1})=\psi^{\prime}(t_{1})-\psi^{\prime}(t_{0})=(t_{1}-t_{0})\psi^{\prime\prime}(\xi)\geq(t_{1}-t_{0})\psi^{\prime\prime}(t_{1}).$
We choose $t_{1}=2t_{0}$, in which case
$\psi^{\prime\prime}(t_{1})=\theta q(1-q)t_{1}^{q-2}=\theta q(1-q)2^{q-2}t_{0}^{q-2}=(1-q)2^{q-2}(q\theta)^{-\frac{1}{1-q}}\sigma^{\frac{2-q}{1-q}},$
and hence
$\psi^{\prime}(t_{1})\geq(t_{1}-t_{0})\psi^{\prime\prime}(t_{1})=t_{0}\psi^{\prime\prime}(t_{1})=(1-q)2^{q-2}\sigma\geq\frac{1}{4}\,(1-q)\sigma.$
Since $\psi(t_{1})>\psi(t_{0})$, it now follows from Lemma 3.2 that
(3.9) $\int_{0}^{+\infty}\exp(\theta t^{q}-\sigma t)\,\mathrm{d}t\leq\mathrm{e}^{-\psi(t_{0})}\Bigg(\sqrt{\frac{2\pi}{\psi^{\prime\prime}(t_{1})}}+\frac{1}{\psi^{\prime}(t_{1})}\Bigg)\\ \leq\frac{1}{\sigma}\bigg(2^{1-\frac{q}{2}}(q\theta)^{\frac{1}{2(1-q)}}\sqrt{\frac{2\pi}{1-q}}\,\sigma^{-\frac{q}{2(1-q)}}+\frac{4}{1-q}\bigg)\exp\Big((1-q)\theta(q\theta/\sigma)^{\frac{q}{1-q}}\Big),$
which amounts to the claimed estimate of the lemma. ∎
4. The instance $0<q<\frac{1}{2}$ connected with analytic non-quasianalyticity
4.1. Background on quasianalyticity and analytic quasianalyticity
Quasianalytic classes were à la mode in the 1920s, with the striking insights of Denjoy [13] and Carleman [11].
If a function on the real line is real-analytic, then, if all its Taylor coefficients vanish at a single point, the function then must vanish identically. This uniqueness property extends to more general $C^{\infty}$-smooth functions on the line, provided we have suitable control on the growth the higher order derivatives in a (locally) uniform sense, as given by a weight sequence. The precise result is known as the Denjoy-Carleman theorem, in which the following dichotomy appears. Either the class is quasianalytic, in the sense that if all the Taylor coefficients of the function vanish at a point, then the function must vanish identically, or the class is non-quasianalytic, meaning that the class contains nontrivial functions with bounded support. As a side note, we observe that Paul Cohen found an elementary proof of the Denjoy-Carleman theorem [12]. Moreover, it might be relevant to mention here that when the uniform control is replaced by $L^{p}$-control, an additional dichotomy shows up in the low exponent regime $0<p<1$ [19].
The concept of analytic quasianalyticity is analogous. If we have a function which is $C^{\infty}$-smooth on the real line, and also possesses a (necessarily unique) holomorphic extension to the upper half-plane, we may still ask when the vanishing of all the Taylor coefficients at a given real point forces the function to vanish everywhere. The natural condition on the (locally) uniform control on the growth of the higher order derivatives remains in force, and the question is asked in terms of the weight sequence controlling the growth. Naturally, the one-sided holomorphic extension property means that less stringent growth control is needed to force the function to vanish everywhere. And when we are in the opposite world of analytic non-quasianalytic classes, that means that that there exist nontrivial functions in the class that are flat at a given real point and holomorphic in the upper half-plane. How should such functions look like? We cannot as in the classical non-quasianalytic case just take such a function and replace it by the 0 function on one side of the given point, as that would ruin the one-sided holomorphic extension property. Typically, if such flat functions exist, there also exist such flat functions with a decay at the given flat point which from the positive imaginary direction is like that of an atomic inner function in the inner-outer factorization theory of the upper half-plane. In [5], Beurling studied weighted spaces on semigroups where analytic non-quasianalyticity is a natural dividing property after application of the Fourier-Laplace transform. Dually, we have instead a space of functions that are defined and holomorphic in the upper half-plane but need not have well-defined boundary values on the real line pointwise. Instead, the functions meet a growth condition in the upper half-plane. Since basically Fourier analysis on the half-line amounts to expansion of the functions in terms of the singular atomic inner functions at the point at infinity, it becomes natural to ask when this carries over to the setting of a finite point in a stable fashion. This stability would require the various atomic inner functions at the given point $x_{0}\in\mathbb{R}$, naturally indexed by a parameter $\alpha\in\mathbb{R}_{\geq 0}$, to generate invariant subspaces $I_{\alpha}(x_{0})$ which get smaller as $\alpha$ increases, and the containments should be strict. This issue appeared later in the work of Korenblum [22] (also in collaboration with Hayman [15]). It was motivated by Korenblum’s earlier work [20], [21] which extended as far as possible the classical Nevanlinna theory as well as the Beurling invariant subspace theorem to the soft topology slow growth spaces $\mathcal{A}^{-\infty}$ consisting of entire functions of finite exponential type in the context of the Poincaré model of the hyperbolic plane. What we will need here is basically the property that the limit of the invariant subspaces $I_{\alpha}(x_{0})$ as $\alpha\to+\infty$ is the trivial subspace $\{0\}$, interpreted in the sense of intersection.
4.2. A boundary decay uniqueness theorem
The following is a prototypical radial boundary decay uniqueness theorem.
Theorem 4.1.
Suppose $F$ is a holomorphic function in the open unit disk $\mathbb{D}$, and that $F$ is bounded. Suppose, moreover, that we have the radial limit
$\lim_{r\to 1^{-}}(1-r)\log|F(r)|=-\infty.$
Then $F$ must vanish identically in $\mathbb{D}$.
This theorem is an immediate consequence of the classical Nevanlinna factorization theory, and we refrain from supplying a proof. We could mention, however, that for Blaschke products
$B(z)=z^{N}\prod_{j}\frac{|a_{j}|}{a_{j}}\frac{a_{j}-z}{1-\bar{a}_{j}z},$
where $N\in\mathbb{Z}_{\geq 0}$ and the sequence $\{a_{j}\}_{j}$ from $\mathbb{D}$ meets the Blaschke condition
$\sum_{j}(1-|z_{j}|)<+\infty,$
it is not hard to establish that
$\limsup_{r\to 1^{-}}(1-r)\log|B(r)|=0,$
so that the only source of a nontrivial limit in this expression could only be the nonzero factors (outer and singular inner). As it happens, only the singular (in fact atomic) inner part can contribute. In fact, the examples of atomic inner functions
$F(z)=\exp\bigg(-\alpha\frac{1+z}{1-z}\bigg)$
for $\alpha\geq 0$ show that any finite limit
$\lim_{r\to 1^{-}}(1-r)\log|F(r)|=-2\alpha$
can be achieved for a nontrivial bounded analytic function. Note that we can use a conformal transformation (or, more generally, a holomorphic transformation), to achieve a weakening of the assumptions of the theorem. This idea is one of the basic ingredients which helped Korenblum developed his non-Nevanlinna theory in [20], [21]. Here, the more precise setting is in part taken from the paper of Hayman and Korenblum [14], [15].
Theorem 4.2.
Suppose $F$ is a holomorphic function in the open unit disk $\mathbb{D}$, and that there exists a holomorphic mapping $\phi:\mathbb{D}\to\mathcal{U}$, where $\mathcal{U}\subset\mathbb{D}$ is a subdomain, and a nontrivial bounded holomorphic function $G$ in $\mathcal{U}$ such that the product $FG$ is bounded on $\mathcal{U}$ as well. The mapping $\phi$ is assumed real-valued on $\mathbb{D}\cap\mathbb{R}$ and strictly increasing on some interval $r_{0},1[$ with $0<r_{0}<1$ and $\lim\varphi(r)=1$ as $r\to 1^{-}$, while
$1-\phi(r)\leq C_{0}(1-r),\qquad r_{0}\leq r<1,$
holds for some positive constant $C_{0}$. Suppose, moreover, that we have the radial limit
$\lim_{r\to 1^{-}}(1-r)\log|F(r)|=-\infty.$
Then $F$ must vanish identically in $\mathbb{D}$.
Proof.
In view of the given decay assumption, we know that
(4.1) $\lim_{r\to 1^{-}}(1-\phi(r))\log|F\circ\phi(r)|=-\infty.$
Moreover, by the inequality $1-\phi(r)\leq C_{0}(1-r)$, it follows from this that
(4.2) $\lim_{r\to 1^{-}}(1-r)\log|F\circ\phi(r)|=-\infty,$
and consequently that
(4.3) $\lim_{r\to 1^{-}}(1-r)\log|(FG)\circ\phi(r)|=$-$\infty$
as well. But now we may apply Theorem 4.1 to the bounded analytic function $(FG)\circ\phi$ in place of $F$ and conclude that $(FG)\circ\phi=0$ identically. But then $FG=0$ on $\mathcal{U}$ and hence $F=0$ identically. ∎
4.3. Application to the setting of the right half-plane and the proof of Theorem 1.17
We need to apply Theorem 4.2 in the setting of the right-half plane rather than the unit disk. In principle, we can use the standard Cayley mapping to obtain the required statement from Theorem 4.2.
Corollary 4.3.
Suppose $F$ is a holomorphic function in the right half-plane $\mathbb{C}_{\mathrm{Re}\,>0}$, and that there exists a holomorphic mapping $\phi:\mathbb{D}\to\mathcal{U}$, where $\mathcal{U}\subset\mathbb{C}_{\mathrm{Re}\,>0}$ is a subdomain, and a nontrivial bounded holomorphic function $G$ on $\mathcal{U}$ such that $FG$ is bounded on $\mathcal{U}$ as well. The mapping $\phi$ is real-valued on $\mathbb{D}\cap\mathbb{R}$ and strictly decreasing on some interval $0,x_{0}[$ with $0<x_{0}<1$ and $\lim\varphi(r)=0$ as $r\to 1^{-}$, while
$1-\phi(x)\leq C_{0}x,\qquad 0<x\leq x_{0},$
holds for some positive constant $C_{0}$. Suppose, moreover, that we have the radial limit
$\lim_{x\to 0^{+}}x\log|F(x)|=-\infty.$
Then $F$ must vanish identically in $\mathbb{C}_{\mathrm{Re}\,>0}$.
Proof of Theorem 1.17.
As in the proof of Theorem 1.14, we may assume that $u=u_{\rm{horz}}=\mathbf{R}[f,0]$ is a unilateral horizontal wave solution, with $f(x)=u(x,0)$ for $x\in\mathbb{R}_{\geq 0}$, so that $f$ is continuous and $f(0)=0$. The given growth control says that
$|u(x,y)|\leq C(y)\,\exp(\theta\,x^{q}),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathcal{Y},$
where $C(y)>0$ may depend arbitrarily on $y\in\mathcal{Y}$. Since $0<q<\frac{1}{2}$ is assumed, the Laplace transform of $u(\cdot,y)$ is holomorphic in the right half-plane $\mathbb{C}_{\mathrm{Re}\,>0}$ for $y\in\mathcal{Y}$, and, in view of Lemma 3.1, we have the estimate, for $\sigma=\mathrm{Re}\,\zeta>0$ and $y\in\mathcal{Y}$:
(4.4) $|\mathbf{L}[u(\cdot,y)](\zeta)|\leq\int_{0}^{+\infty}|u(x,y)|\,\mathrm{e}^{-x\mathrm{Re}\,\zeta}\mathrm{d}x\leq C(y)\int_{0}^{+\infty}\exp(\theta x^{q}-\sigma x)\mathrm{d}x\\ \leq C(y)\,\frac{1}{\sigma}\big(A(q,\theta)\sigma^{-\frac{q}{2(1-q)}}+B_{q}\big)\exp\big(D(q,\theta)\,\sigma^{-\frac{q}{1-q}}\big),$
where the constants are as in the lemma. Next, in view of (2.6), we know that
$\mathbf{L}[u(\cdot,y)](\zeta)=\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta),$
so that (4.4) entails that
(4.5) $|\mathbf{L}[f](\zeta)|\leq C(y)\,\frac{1}{\sigma}\big(A(q,\theta)\sigma^{-\frac{q}{2(1-q)}}+B_{q}\big)\\ \times\exp\big(D(q,\theta)\,\sigma^{-\frac{q}{1-q}}+y\mathrm{Re}\,(1/\zeta)\big),\qquad y\in\mathcal{Y}.$
It now follows that for $\zeta=\sigma>0$, we have that
(4.6) $\sigma\log|\mathbf{L}[f](\sigma)|\leq\sigma\log\frac{C(y)}{\sigma}\\ +\sigma\log\big(A(q,\theta)\sigma^{-\frac{q}{2(1-q)}}+B_{q}\big)+D(q,\theta)\,\sigma^{\frac{1-2q}{1-q}}+y,\qquad y\in\mathcal{Y}.$
We let $\sigma\to 0^{+}$, and observe that since $0<q<\frac{1}{2}$, it follows from (4.6) that
(4.7) $\limsup_{\sigma\to 0^{+}}\sigma\log|\mathbf{L}[f](\sigma)|\leq y,\qquad y\in\mathcal{Y},$
and since $y\in\mathcal{Y}$ is arbitrary, and $\inf\mathcal{Y}=$-$\infty$ holds, it then follows that
(4.8) $\lim_{\sigma\to 0^{+}}\sigma\log|\mathbf{L}[f](\sigma)|=-\infty.$
The construction of the domain $\mathcal{U}$ and associated mapping $\phi:\mathbb{D}\to\mathcal{U}$ is somewhat complicated but carried out by Hayman and Korenblum in [14], [15] in the context of the condition
(4.9) $\int_{0}^{1}\sqrt{\frac{\kappa(t)}{t}}\mathrm{d}t<+\infty,$
where the function $\kappa$ measures the growth. In our situation, $\kappa$ is, if we neglect logarithmic contributions,
$\kappa(t)=D(q,\theta)\,t^{-\frac{q}{1-q}},$
and since $0<q<\frac{1}{2}$, the condition (4.9) is automatically met. By the work of Hayman and Korenblum, then, we are then in the setting of Corollary 4.3, and conclude that $F=\mathbf{L}[f]$ must vanish identically, and hence $f=0$ almost everywhere on $\mathbb{R}_{\geq 0}$, so that $u=\mathbf{R}[f,0]=\mathbf{R}[0,0]=0$ in the quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, as claimed. ∎
Remark 4.4.
The condition (4.9) is understood as a version of the criterion for ananlytic non-quasianalyticity.
5. Uniqueness in the skewed exponential growth case $\frac{1}{2}<q<1$
5.1. Skewed exponential growth
As our starting point, we have a solution $u$ to the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the spacelike quarter-plane $\mathbb{R}_{>0}\times\mathbb{R}_{<0}$ which extends continuously to the closed quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, with vanishing vertical boundary values:
$u(0,y)=0,\qquad y\in\mathbb{R}_{\leq 0}.$
As explained in the proof of Theorem 1.14, $u=\mathbf{R}[f,0]$ if the continuous function $f$ is given by $f(x)=u(x,0)$. We assume that $\frac{1}{2}<q<1$, and that
(5.1) $|u(x,y)|\leq C_{0}\exp(\theta_{1}|x|^{q}+\theta_{2}|y|^{q^{\prime\prime}}),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathcal{Y},$
holds for some positive constants $C_{0}$, and $\theta_{1},\theta_{2}$, where $q^{\prime\prime}=q/(2q-1)$. Here, $\mathcal{Y}$ is a subset of the half-line $\mathbb{R}_{\leq 0}$ with the properties prescribed by Theorem 1.20. We apply the Laplace transform with respect to the $x$ variable:
$\mathbf{L}[u(\cdot,y)](\zeta)=\int_{0}^{+\infty}u(x,y)\,\mathrm{e}^{-x\zeta}\mathrm{d}x,\qquad\mathrm{Re}\,\zeta>0,$
which defines a holomorphic function in the indicated half-plane for each $y\in\mathbb{R}_{\leq 0}$. In view of (2.6),
$\mathbf{L}[u(\cdot,y)](\zeta)=\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta),$
and by the estimate (5.1), we have, for $\sigma=\mathrm{Re}\,\zeta>0$,
(5.2) $|\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta)|=\mathbf{L}[u(\cdot,y)](\zeta)|\leq\int_{0}^{+\infty}|u(x,y)|\,\mathrm{e}^{-\sigma x}\mathrm{d}x\\ \leq C_{0}\exp(\theta_{2}|y|^{q^{\prime\prime}})\int_{0}^{+\infty}\exp(\theta_{1}x^{q}-\sigma x)\,\mathrm{d}x,\qquad y\in\mathcal{Y},$
so that
(5.3) $|\mathbf{L}[f](\zeta)|\leq C_{0}\exp\big(\theta_{2}|y|^{q^{\prime\prime}}+\mathrm{Re}\,(y/\zeta)\big)\int_{0}^{+\infty}\exp(\theta_{1}x^{q}-\sigma x)\,\mathrm{d}x,\qquad y\in\mathcal{Y},$
where again we write $\sigma=\mathrm{Re}\,\zeta>0$. Here, we may optimize over $y\in\mathcal{Y}$, which yields
(5.4) $|\mathbf{L}[f](\zeta)|\leq C_{0}\exp\Big(\inf_{y\in\mathcal{Y}}\big\{\theta_{2}|y|^{q^{\prime\prime}}+\mathrm{Re}\,(y/\zeta)\big\}\Big)\int_{0}^{+\infty}\exp(\theta_{1}x^{q}-\sigma x)\,\mathrm{d}x.$
If we write
$\psi(y)=\theta_{2}|y|^{q^{\prime\prime}}-|y|\,\mathrm{Re}\,(1/\zeta),$
we may first minimize over all $y\in\mathbb{R}_{\leq 0}$ and then hope that the minimum over $y\in\mathcal{Y}$ is not too much bigger. The minimum over $y\in\mathbb{R}_{\leq 0}$ is attained at the point
(5.5) $y_{\star}(\zeta,\theta_{2})=-\bigg(\frac{\mathrm{Re}\,(1/\zeta)}{\theta_{2}q^{\prime\prime}}\bigg)^{\frac{1}{q^{\prime\prime}-1}},$
where we suppress indication of the variable $q$ as it is kept fixed, and then
$\psi(y_{\star}(\zeta,\theta_{2}))=-E(q,\theta_{2})(\mathrm{Re}\,(1/\zeta))^{\frac{q}{1-q}},$
where
$E(q,\theta_{2}):=\bigg(\frac{1}{q}-1\bigg)(q^{\prime\prime}\theta_{2})^{1-\frac{q}{1-q}}.$
At some other point $y\in\mathbb{R}_{\leq 0}$, we may apply a version of Taylor’s formula:
(5.6) $\psi(y)=\psi(y_{\star}(\zeta,\theta_{2}))+(y-y_{\star}(\zeta,\theta_{2}))\psi^{\prime}(y_{\star}(\zeta,\theta_{2}))+\frac{1}{2}(y-y_{\star}(\zeta,\theta_{2})^{2}\psi^{\prime\prime}(\xi)\\ =\psi(y_{\star}(\zeta,\theta_{2}))+\frac{1}{2}(y-y_{\star}(\zeta,\theta_{2}))^{2}\psi^{\prime\prime}(\xi),$
since $\psi(y_{\star}(\zeta,\theta_{2}))=0$ must hold at the minimum. Here, $\xi$ stands for an intermediate point between $y_{\star}(\zeta,\theta_{2})$ and $y$. Since
$\psi^{\prime\prime}(y)=\theta_{2}q^{\prime\prime}(q^{\prime\prime}-1)|y|^{q^{\prime\prime}-2},$
it now follows that
(5.7) $\psi(y)=\psi(y_{\star}(\zeta,\theta_{2}))+\frac{1}{2}(y-y_{\star}(\zeta,\theta_{2}))^{2}\psi^{\prime\prime}(\xi)\\ =-E(q,\theta_{2})(\mathrm{Re}\,(1/\zeta))^{\frac{q}{1-q}}+\frac{1}{2}(y-y_{\star}(\zeta,\theta_{2}))^{2}\theta_{2}q^{\prime\prime}(q^{\prime\prime}-1)|\xi|^{q^{\prime\prime}-2}.$
Lemma 5.1.
Suppose $\mathcal{Y}\subset\mathbb{R}_{\leq 0}$ and the point $y_{1}\in\mathbb{R}_{\leq 0}$ are as in the formulation of Theorem 1.20. If $y_{2}:=\min\{-1,y_{1},-(2M)^{2/q^{\prime\prime}}\}$, and if $y_{\star}<(M+1)y_{2}$, then there exists a point $y\in\mathcal{Y}$ with $y<y_{2}$ such that
$y-M|y|^{1-\frac{q^{\prime\prime}}{2}}\leq y_{\star}\leq y+M|y|^{1-\frac{q^{\prime\prime}}{2}},$
and hence
$2y_{\star}\leq y\leq\frac{2}{3}y_{\star}.$
Proof.
If $y\in\mathcal{Y}$ with $y_{2}\leq y\leq 0$, we use the inequality $0\leq R(y)\leq|y|$ to show that we have the containment of intervals
$y-MR(y),y+MR(y)]\subset[(M+1)y_{2},(1-M)y_{2}],$
so that in view of the assumed inequality $y_{\star}<(M+1)y_{2}$, it follows that
$y_{\star}\notin[y-MR(y),y+MR(y)],\qquad y\in\mathcal{Y},\,\,\,y_{2}\leq y\leq 0,$
so that in view of the asymptotic $q$-covering property, there must instead exist a $y\in\mathcal{Y}$ with $y<y_{2}$ such that
$y_{\star}\in[y-MR(y),y+MR(y)],$
and since $y<y_{2}\leq-1$, we know that $R(y)=|y|^{1-\frac{q^{\prime\prime}}{2}}$, and hence
(5.8) $y(1+M|y|^{-\frac{q^{\prime\prime}}{2}})=y-M|y|^{1-\frac{q^{\prime\prime}}{2}}\leq y_{\star}\leq y+M|y|^{1-\frac{q^{\prime\prime}}{2}}=y(1-M|y|^{-\frac{q^{\prime\prime}}{2}}).$
Moreover, since $y<y_{2}\leq-(2M)^{2/q^{\prime\prime}}$, it follows that $|y|^{-q^{\prime\prime}/2}<(2M)^{-1}$, and hence, by (5.8),
(5.9) $\frac{3}{2}y\leq y_{\star}\leq\frac{1}{2}y,$
which is equivalent to
(5.10) $2y_{\star}\leq y\leq\frac{2}{3}y_{\star},$
as claimed. ∎
Lemma 5.2.
Suppose $y,y_{\star}\in\mathbb{R}_{<0}$ are such that $y\leq-1$ and
$y-M|y|^{1-\frac{q^{\prime\prime}}{2}}\leq y_{\star}\leq y+M|y|^{1-\frac{q^{\prime\prime}}{2}},$
and
$2y_{\star}\leq y\leq\frac{2}{3}y_{\star}.$
Then, if $\xi$ is an intermediate point between $y_{\star}$ and $y$, it follows that
$0\leq(y-y_{\star})^{2}|\xi|^{q^{\prime\prime}-2}\leq 2^{|q^{\prime\prime}-2|}M^{2}.$
Proof.
Since
$(y-y_{\star})^{2}\leq M^{2}|y|^{2-q^{\prime\prime}},$
and, in addition, we check that
$|\xi|^{q^{\prime\prime}-2}\leq 2^{|q^{\prime\prime}-2|}|y|^{q^{\prime\prime}-2},$
so that, consequently,
$0\leq(y-y_{\star})^{2}|\xi|^{q^{\prime\prime}-2}\leq 2^{|q^{\prime\prime}-2|}M^{2},$
as claimed. ∎
Proof of Theorem 1.20.
We let $y_{\star}(\zeta,\theta_{2})$ be given by (5.5), and assume that $y_{\star}(\zeta,\theta_{2})<(M+1)y_{2}$, where $y_{2}$ is as in Lemma 5.1. Then, according to Lemma 5.1, there exists a point $y\in\mathcal{Y}$ with $y<y_{2}$ such that in accordance with Lemma 5.2 combined with equation (5.7), we have
(5.11) $\psi(y)\leq-E(q,\theta_{2})(\mathrm{Re}\,(1/\zeta))^{\frac{q}{1-q}}+2^{|q^{\prime\prime}-2|-1}\theta_{2}q^{\prime\prime}(q^{\prime\prime}-1)M^{2}.$
As we insert this point into the estimate (5.4), we find that
(5.12) $|\mathbf{L}[f](\zeta)|\leq C_{1}\exp\Big(-E(q,\theta_{2})(\mathrm{Re}\,(1/\zeta))^{\frac{q}{1-q}}\Big)\int_{0}^{+\infty}\exp(\theta_{1}x^{q}-\sigma x)\,\mathrm{d}x,\qquad\zeta\in\Omega_{2},$
where
$C_{1}:=C_{0}\exp\big(2^{|q^{\prime\prime}-2|-1}\theta_{2}q^{\prime\prime}(q^{\prime\prime}-1)M^{2}\big),$
and $\Omega_{2}$ stands for the domain induced by the requirement that $y_{\star}(\zeta,\theta_{2})<y_{2}$, that is,
(5.13) $\Omega_{2}:=\Big\{z\in\mathbb{C}_{\mathrm{Re}\,>0}:\,\,\mathrm{Re}\,(1/\zeta)>\theta_{2}q^{\prime\prime}(M+1)^{q^{\prime\prime}-1}|y_{2}|^{q^{\prime\prime}-1}\Big\},$
which is a horocyclic disk tangential to the imaginary line at the origin. We may now apply Lemma 3.1 to estimate the integral in (5.12), which gives, for $\sigma\in\mathbb{R}_{>0}\cap\Omega_{2}$,
(5.14) $|\mathbf{L}[f](\sigma)|\leq\frac{C_{1}}{\sigma}\big(A(q,\theta_{1})\sigma^{-\frac{q}{2(1-q)}}+B(q)\big)\exp\Big((D(q,\theta_{1})-E(q,\theta_{2}))\,\sigma^{-\frac{q}{1-q}}\Big).$
while more generally, we have
(5.15) $|\mathbf{L}[f](\zeta)|\leq\frac{C_{1}}{\sigma}\big(A(q,\theta_{1})\sigma^{-\frac{q}{2(1-q)}}+B(q)\big)\\ \times\exp\Big((D(q,\theta_{1}))\,\sigma^{-\frac{q}{1-q}}-E(q,\theta_{2})(\mathrm{Re}\,(1/\zeta))^{\frac{q}{1-q}}\Big),$
for $\zeta\in\Omega_{2}$. Since
$\mathrm{Re}\,\frac{1}{\zeta}=\frac{\mathrm{Re}\,\zeta}{|\zeta|^{2}}=\frac{\sigma}{|\zeta|^{2}}=\frac{1}{\sigma}\,\bigg(\frac{\sigma}{|\zeta|}\bigg)^{2},$
it follows from (5.15) that
(5.16) $|\mathbf{L}[f](\zeta)|\leq\frac{C_{1}}{\sigma}\big(A(q,\theta_{1})\sigma^{-\frac{q}{2(1-q)}}+B(q)\big),\qquad\zeta\in\Omega_{2}\cap\Omega_{3},$
holds, where $\Omega_{3}$ denotes the domain in the half-plane $\mathbb{C}_{\mathrm{Re}\,>0}$ where
(5.17) $D(q,\theta_{1})<E(q,\theta_{2})\,\bigg(\frac{\sigma}{|\zeta|}\bigg)^{\frac{2q}{1-q}},$
that is, where
(5.18) $\frac{\sigma^{2}}{|\zeta|^{2}}>\bigg(\frac{D(q,\theta_{1})}{E(q,\theta_{2})}\bigg)^{\frac{1}{q}-1}={(q\theta_{1})^{\frac{1}{q}}}{(q^{\prime\prime}\theta_{2})^{\frac{1}{q^{\prime\prime}}}},$
which is an angular opening with apex at the origin. By assumption, we have that
${(q\theta_{1})^{\frac{1}{q}}}{(q^{\prime\prime}\theta_{2})^{\frac{1}{q^{\prime\prime}}}}<\cos^{2}\frac{\pi(1-q)}{2q}<1,$
so that $D(q,\theta_{1})<E(q,\theta_{2})$ and hence we have exponential decay at the given rate along the real line as $\sigma\to 0^{+}$ in (5.14). If we write
${(q\theta_{1})^{\frac{1}{q}}}{(q^{\prime\prime}\theta_{2})^{\frac{1}{q^{\prime\prime}}}}=\cos^{2}\frac{\pi\alpha}{2},\qquad\frac{1-q}{q}<\alpha<1,$
then $\alpha=\alpha(q,\theta_{1},\theta_{2})$, and the region $\Omega_{3}$ is an angular opening toward the right, centered at the origin, symmetric about the real line, with aperture $\pi\alpha$. Under the change-of-variables $\zeta=\eta^{\alpha}$, then $\zeta\in\Omega_{3}$ if and only if $\eta\in\mathbb{C}_{\mathrm{Re}\,>0}$. Here, the power is defined based on the principal branch of the logarithm. We put
$F(\eta):=\mathbf{L}[f](\eta^{\alpha}),$
which is then holomorphic in the symmetric angular opening of aperture $\pi/\alpha$, and hence in particular in the right half-plane $\mathbb{C}_{\mathrm{Re}\,>0}$. It follows from the estimate (5.16) that uniformly,
(5.19) $|F(\eta)|=\mathrm{O}\big(|\eta|^{-{\alpha}}\big(|\eta|^{-\frac{\alpha q}{2(1-q)}}+1)\big)\big),\qquad\eta\in\mathbb{C}_{\mathrm{Re}\,>0}\cap\tilde{\Omega}_{2},$
provided that $\eta\in\tilde{\Omega}_{2}$ means that $\zeta=\eta^{\alpha}\in\Omega_{2}$. At the same time, we have exponential decay in accordance with (5.14):
(5.20) $|F(\rho)|=\mathrm{O}\big(\rho^{-\alpha}\big(\rho^{-\frac{\alpha q}{2(1-q)}}+1\big)\exp\Big((D(q,\theta_{1})-E(q,\theta_{2}))\,\rho^{-\frac{\alpha q}{1-q}}\Big)\Big),$
uniformly over $\rho\in\mathbb{R}_{>0}$. Since $D(q,\theta_{1})<E(q,\theta_{2})$ and $\alpha q/(1-q)>1$, it is an immediate consequence of (5.20) that
(5.21) $\lim_{\rho\to 0^{+}}\rho\log|F(\rho)|=-\infty.$
If $F$ were bounded in the half-plane $\mathbb{C}_{\mathrm{Re}\,>0}$, we could then immediately invoke Theorem 4.1 and obtain that $F=0$ holds identically. But all we have is the estimate (5.19), which allows for slight growth at the boundary point 0. It is now easy to supply an outer function $G$ in the domain $\mathbb{C}_{\mathrm{Re}\,>0}\cap\tilde{\Omega}_{2}$ whose modulus of the boundary values equal those of the right-hand side estimate:
$|G(\eta)|=\min\big\{1,|\eta|^{\alpha},|\eta|^{\alpha+\frac{\alpha q}{2(1-q)}}\big\},\qquad\eta\in\partial(\mathbb{C}_{\mathrm{Re}\,>0}\cap\tilde{\Omega}_{2}),$
in which case $G$ is nontrivial and bounded, and more importantly, the product $FG$ is bounded too, and has
(5.22) $\lim_{\rho\to 0^{+}}\rho\log|(FG)(\rho)|=$-$\infty$
as a consequence of (5.21). Now the geometry of the simply connected domain $\mathbb{C}_{\mathrm{Re}\,>0}\cap\tilde{\Omega}_{2}$ is such that it looks like a half-plane near the origin, and then, the conformal map $\phi:\mathbb{D}\to\mathbb{C}_{\mathrm{Re}\,>0}\cap\tilde{\Omega}_{2}$ which maps $\phi(1)=0$ extends conformally across the point $1$, so that the conditions of Theorem 4.1 are all met, and we conclude that $FG=0$. Since $G$ is nontrivial, it follows that $F=0$ identically on $\mathbb{C}_{\mathrm{Re}\,>0}$, which gives that $\mathbf{L}[f]=0$ and hence $f=0$. This completes the proof of the theorem. ∎
6. Uniqueness in the critical growth rate case $q=\frac{1}{2}$
6.1. The setting of critical growth $q=\frac{1}{2}$
As our starting point, we have a solution to the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the spacelike quarter-plane $\mathbb{R}_{>0}\times\mathbb{R}_{<0}$ which extends continuously to the closed quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$, with vanishing vertical boundary values:
$u(0,y)=0,\qquad y\in\mathbb{R}_{\leq 0}.$
As explained in the proof of Theorem 1.14, $u=\mathbf{R}[f,0]$ if the continuous function $f$ is given by $f(x)=u(x,0)$. We consider $q=\frac{1}{2}$, and suppose that the growth of $u$ is controlled by
(6.1) $|u(x,y)|\leq C_{0}\exp\big(\theta_{1}x^{\frac{1}{2}}+\beta(y)),\qquad(x,y)\in\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0},$
where $\beta(y)$ is the function
$\beta(y)=|y|\exp(\theta_{2}\sqrt{|y|}),$
which even, and on $\mathbb{R}_{\geq 0}$ it is strictly increasing and strictly convex, by inspection of the first and second derivatives.
Proof of Theorem 1.22.
Since we easily establish that
$\exp(\theta_{2}\sqrt{|y|})\leq\mathrm{e}^{\theta_{2}}+|y|\exp(\theta_{2}\sqrt{|y|})=\beta(y)+\mathrm{O}(1),\qquad y\in\mathbb{R},$
the assumption of the theorem entails that we are in the setting of the inequality (6.1). We then apply the Fourier-Laplace transform with respect to $x$:
$\mathbf{L}[u(\cdot,y)](\zeta)=\int_{0}^{+\infty}\mathrm{e}^{-x\zeta}u(x,y)\,\mathrm{d}x,\qquad\zeta\in\mathbb{C}_{\mathrm{Re}\,>0},$
which defines a holomorphic function in the given half-plane. Moreover, in view of the bound (6.1), we have
(6.2) $|\mathbf{L}[u(\cdot,y)](\zeta)|\leq C_{0}\exp(\beta(y))\int_{0}^{+\infty}\mathrm{e}^{\theta_{1}x^{\frac{1}{2}}-x\mathrm{Re}\,\zeta}\,\mathrm{d}x,\qquad(\zeta,y)\in\mathbb{C}_{\mathrm{Re}\,>0}\times\\ R_{\leq 0},$
and since
$\mathbf{L}[u(\cdot,y)](\zeta)=\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta),$
we derive that
(6.3) $|\mathbf{L}[f](\zeta)|\leq C_{0}\exp\big(\beta(y)+y\mathrm{Re}\,(1/\zeta)\big)\int_{0}^{+\infty}\mathrm{e}^{\theta_{1}x^{\frac{1}{2}}-x\mathrm{Re}\,\zeta}\,\mathrm{d}x,\qquad\zeta\in\mathbb{C}_{\mathrm{Re}\,>0},$
where $y\in\mathbb{R}_{\leq 0}$ is arbitrary. If we write
$\beta^{\ast}(t):=\sup\{yt-\beta(y):\,\,\,y\in\mathbb{R}_{\geq 0}\}\geq 0,$
the result is another convex function that is, in the sense of Legendre, the convex conjugate to $\beta$ itself. By performing asymptotic calculations, we can establish that
(6.4) $\beta^{\ast}(t)=\frac{1}{\theta_{2}^{2}}\,t\log^{2}t-\frac{2}{\theta_{2}^{2}}\,t\log t\log\log t+\mathrm{O}(t\log t)$
holds asymptotically as $t\to+\infty$. It now follows from (6.3) that
(6.5) $|\mathbf{L}[f](\zeta)|\leq C_{0}\exp\big(-\beta^{\ast}(\mathrm{Re}\,(1/\zeta))\big)\int_{0}^{+\infty}\mathrm{e}^{\theta_{1}x^{\frac{1}{2}}-x\mathrm{Re}\,\zeta}\,\mathrm{d}x,\qquad\zeta\in\mathbb{C}_{\mathrm{Re}\,>0},$
and if we use Lemma 3.1 to estimate the integral, we find that
$\int_{0}^{+\infty}\mathrm{e}^{\theta_{1}x^{\frac{1}{2}}-x\sigma}\,\mathrm{d}x\leq\frac{1}{\sigma}\big(2^{3/4}\sqrt{\pi}\theta_{1}\sigma^{-\frac{1}{2}}+8\big)\exp\bigg(\frac{\theta_{1}^{2}}{4\sigma}\bigg),$
so that if we write $\mathrm{Re}\,\zeta=\sigma$, then
(6.6) $|\mathbf{L}[f](\zeta)|\leq C_{0}\frac{1}{\sigma}\big(2^{3/4}\sqrt{\pi}\theta_{1}\sigma^{-\frac{1}{2}}+8\big)\exp\bigg(-\beta^{\ast}(\mathrm{Re}\,(1/\zeta))+\frac{\theta_{1}^{2}}{4\sigma}\bigg),$
for $\mathrm{Re}\,\zeta=\sigma>0$. Since the point we are interested in is actually the origin $\zeta=0$, which after inversion becomes the point at infinity, it makes sense to consider the function
$F(\zeta):=\mathbf{L}[f](1/\zeta),\qquad\zeta\in\mathbb{C}_{>0},$
for which the estimate (6.6) becomes
(6.7) $|F(\zeta)|\leq C_{0}\frac{|\zeta|^{2}}{\sigma}\big(2^{3/4}\sqrt{\pi}\theta_{1}|\zeta|\sigma^{-\frac{1}{2}}+8\big)\exp\bigg(-\beta^{\ast}(\sigma)+\frac{\theta_{1}^{2}|\zeta|^{2}}{4\sigma}\bigg),\qquad\zeta\in\mathbb{C}_{\mathrm{Re}\,>0}.$
If we consider the region where
$\Omega_{1}:=\bigg\{\zeta\in\mathbb{C}_{\mathrm{Re}\,>0}:\,\,\,\sigma\beta^{\ast}(\sigma)>\frac{\theta_{1}^{2}|\zeta|^{2}}{4}\bigg\},$
then
(6.8) $|F(\zeta)|\leq C_{0}\frac{|\zeta|^{2}}{\sigma}\big(2^{3/4}\sqrt{\pi}\theta_{1}|\zeta|\sigma^{-\frac{1}{2}}+8\big),\qquad\zeta\in\Omega_{1},$
while
(6.9) $|F(\sigma)|\leq C_{0}\big(2^{3/4}\sqrt{\pi}\theta_{1}\sigma^{\frac{3}{2}}+8\sigma\big)\exp\bigg(-\beta^{\ast}(\sigma)+\frac{\theta_{1}^{2}\sigma}{4}\bigg),\qquad 0<\sigma<+\infty.$
We factor $\beta^{\ast}(t)=t\alpha^{\ast}(t)$, and see that the condition defining the set $\Omega_{1}$ is
$\frac{|\zeta|^{2}}{\sigma^{2}}<\frac{4}{\theta_{1}^{2}}\alpha^{\ast}(\sigma),\qquad\zeta\in\Omega_{1},$
that is,
$\frac{\sigma}{|\zeta|}>\frac{\theta_{1}}{2}\,\alpha^{\ast}(\sigma)^{-\frac{1}{2}},\qquad\zeta\in\Omega_{1}.$
In the complement, we have instead
$\frac{\sigma}{|\zeta|}\leq\frac{\theta_{1}}{2}\,\alpha^{\ast}(\sigma)^{-\frac{1}{2}},\qquad\zeta\in\mathbb{C}_{\mathrm{Re}\,>0}\setminus\Omega_{1}.$
Moreover, in view of (6.4), we see that
(6.10) $\alpha^{\ast}(\sigma)^{-\frac{1}{2}}=\frac{\theta_{2}}{\log\sigma}-\theta_{2}\frac{\log\log\sigma}{\log^{2}\sigma}+\mathrm{O}\bigg(\frac{1}{\log^{2}\sigma}\bigg)$
as $\sigma\to+\infty$. Along the boundary $\partial\Omega_{1}$, we have equality,
$\frac{\sigma}{|\zeta|}=\frac{\theta_{1}}{2}\,\alpha^{\ast}(\sigma)^{-\frac{1}{2}}=\frac{\theta_{1}\theta_{2}}{2}\bigg(\frac{1}{\log\sigma}+\frac{\log\log\sigma}{\log^{2}\sigma}+\mathrm{O}(1)\bigg),\qquad\zeta\in\partial\Omega_{1},$
where we used the asymptotics in (6.10), so that
$\rho:=|\zeta|=\frac{2}{\theta_{1}}\sigma\alpha^{\ast}(\sigma)^{\frac{1}{2}}=\frac{2\sigma}{\theta_{1}\theta_{2}}\big(\log\sigma+\log\log\sigma+\mathrm{O}(1)\big),$
as $\sigma\to+\infty$, which relationship between $\sigma$ and $\rho$ we may invert, asymptotically:
$\sigma=\frac{\theta_{1}\theta_{2}}{2}\bigg(\frac{\rho}{\log\rho}+\frac{\rho\log\log\log\rho}{\log^{2}\rho}+\mathrm{O}\bigg(\frac{\rho}{\log^{2}\rho}\bigg)\bigg)=\frac{\theta_{1}\theta_{2}}{2}\frac{\rho}{\log\rho}\big(1+\mathrm{o}(1)\big),$
so that
$\frac{\sigma}{\rho}=\frac{\theta_{1}\theta_{2}}{2\log\rho}\,\big(1+\mathrm{o}(1)\big),$
as $\rho\to+\infty$. This means that the symmetric opening angle of $\Omega_{1}$ on the radius $|\zeta|=\rho$ is $\pi-2\epsilon$, where
(6.11) $\sin\epsilon=\frac{\theta_{1}\theta_{2}}{2\log\rho}\big(1+\mathrm{o}(1)\big),$
and the asymptotics is as $\rho\to+\infty$. Let us consider a slightly smaller domain $\Omega_{2}\subset\Omega_{1}$ which is simply connected and starlike, given as
$\Omega_{2}=\{r\,\mathrm{e}^{\mathrm{i}\varphi}:\,\,\,\,|\varphi|<\tfrac{1}{2}\pi-\epsilon(r)\},$
where $\epsilon:\,\mathbb{R}_{\geq 0}\to[0,\frac{1}{2}\pi]$ is a decreasing function with $\epsilon(r)=\frac{1}{2}\pi$ for $0\leq r\leq r_{0}$ with some $r_{0}>0$, and the asymptotic decay rate
$\epsilon(r)=\frac{\theta_{3}}{2\log r},\qquad R\leq r<+\infty,\qquad\theta_{1}\theta_{2}<\theta_{3}<2\pi.$
Note that this agrees with the asymptotic angle dependence on the radius in equation (6.11), and the containment $\Omega_{2}\subset\Omega_{1}$. Next, suppose $P_{\infty}$ is a positive harmonic function in $\Omega_{2}$ which vanishes along the boundary $\partial\Omega_{2}$. Projectively, there is only one such function, which we understand as meaning that any other such function is the positive constant multiple of the first function we found. We may think of $P_{\infty}$ as the Poisson kernel for the point at infinity, considered as a boundary point of $\Omega_{2}$. The classical Ahlfors-Warschawski theorem [25] characterizes geometrically the growth of the function $P_{\infty}$ along the positive real line. Indeed,
$P_{\infty}(\sigma)\asymp\exp\bigg(\int_{r_{1}}^{\rho}\frac{\pi}{\pi-2\epsilon(r)}\,\frac{\mathrm{d}r}{r}\bigg)$
in the asymptotic sense as $\sigma\to+\infty$, and where the radius $r_{1}$ is selected such that, e.g., $\epsilon(r_{1})=\frac{1}{4}\pi$. We then calculate that if $R\geq r_{1}$ is fixed, we obtain
(6.12) $\int_{R}^{\rho}\frac{\pi}{\pi-2\epsilon(r)}\frac{\mathrm{d}r}{r}=\int_{R}^{\rho}\bigg(1+\frac{2}{\pi}\epsilon(r)+\mathrm{O}(\epsilon(r)^{2})\bigg)\frac{\mathrm{d}r}{r}\\ =\int_{R}^{\rho}\bigg(1+\frac{1}{\pi}\frac{\theta_{3}}{\log r}+\mathrm{O}((\log r)^{-2})\bigg)\frac{\mathrm{d}r}{r}\\ =\log\rho-\log R+\frac{\theta_{3}}{\pi}\big(\log\log\rho-\log\log R\big)+\mathrm{O}(1),$
as $\rho\to+\infty$. Since $R$ is kept fixed, we may replace $r_{1}$ by $R$ in the Ahlfors-Warschawski theorem, so that
(6.13) $P_{\infty}(\sigma)\asymp\exp\bigg(\int_{R}^{\rho}\frac{\pi}{\pi-2\epsilon(r)}\,\frac{\mathrm{d}r}{r}\bigg)\asymp\frac{\rho}{R}\bigg(\frac{\log\rho}{\log R}\bigg)^{\frac{\theta_{3}}{\pi}}$
as $\rho\to+\infty$. We now return to the estimates (6.8) and (6.9). The growth bound (6.8), which automatically holds on $\Omega_{2}$, together with the simple geometry of $\Omega_{2}$, gives that there exists an outer function on $\Omega_{2}$, such that
$|G(\zeta)|=\min\bigg\{\frac{\sigma^{\frac{3}{2}}}{|\zeta|^{3}},\frac{\sigma}{|\zeta|^{2}}\bigg\},\qquad\zeta\in\partial\Omega_{2}.$
By (6.8), we have that
$\big|F(\zeta)G(\zeta)\big|\leq C_{1},\qquad\zeta\in\Omega_{1},$
for some constant $C_{1}$. Moreover, by (6.9), we also have
(6.14) $\big|F(\zeta)G(\zeta)\big|\leq C_{2}\,\exp\bigg(-\beta^{\ast}(\sigma)+\frac{\theta_{1}|\zeta|^{2}}{4\sigma}\bigg),\qquad\zeta\in\Omega_{1},$
for some positive constant $C_{2}$. Since we know that
$\beta^{\ast}(\sigma)=\frac{1}{\theta_{2}^{2}}(1+\mathrm{o}(1))\,\sigma\log^{2}\sigma,$
it now follows from this in combination with (6.13) that for each $\lambda\in\mathbb{R}_{>0}$,
(6.15) $\big|{F(\sigma)}{G(\sigma)}\big|=\mathrm{O}\big(\exp\big(-\lambda P_{\infty}(\sigma)\big)\big)\,\,\,\text{as}\,\,\,\sigma\to+\infty.$
This is basically (after a suitable conformal mapping) the same as the setting of Theorem 4.1, and it follows that $FG=0$ and hence $F=0$ and equivalently $f=0$ on $\mathbb{R}_{\geq}0$. Consequently, $u=\mathbf{R}[f,0]=\mathbf{R}[0,0]=0$ on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$. ∎
7. Appendix on gluing and on Riemann’s solution
7.1. Gluing along characteristic lines
We turn to the proof of Theorem 1.7.
Proof of Theorem 1.7.
The assertion is trivial unless $\Omega\cap L\neq\emptyset$, so we will make this assumption. In view of the stated characterization of continuous solutions to the Klein-Gordon equation in terms of the quadrature identity (1.6) for any rectangle $a,a^{\prime}]\times[b,b^{\prime}]$ contained in $\Omega$, we need to check that
(7.1) $\int_{a}^{a^{\prime}}\int_{b}^{b^{\prime}}u(s,t)\mathrm{d}t\mathrm{d}s=u(a,b^{\prime})+u(a^{\prime},b)-u(a,b)-u(a^{\prime},b^{\prime}).$
By Lorentz group invariance, it is sufficient to handle the instance when $L$ is a horizontal line $u=b^{\prime\prime}$. The distributional interpretation of continuous solutions to the Klein-Gordon equation gives that we know that (7.1) holds for all $a,a^{\prime}]\times[b,b^{\prime}]\subset\Omega\setminus L$, that is, when $a,a^{\prime}]\times[b,b^{\prime}]\subset\Omega$ and $b^{\prime\prime}\notin[b,b^{\prime}]$. In a next step, we may use the continuity of $u$ to allow $b$ or $b^{\prime}$ to approach the value $b^{\prime\prime}$, which establishes (7.1) for $a,a^{\prime}]\times[b,b^{\prime}]\subset\Omega$ and $b^{\prime\prime}\notin]b,b^{\prime}[$. Finally, when $b^{\prime\prime}\in]b,b^{\prime}[$, we split $b,b^{\prime}]=[b,b^{\prime\prime}]\cup[b^{\prime\prime},b^{\prime}]$, and since, by the argument we already carried out,
(7.2) $\int_{a}^{a^{\prime}}\int_{b}^{b^{\prime\prime}}u(s,t)\mathrm{d}t\mathrm{d}s=u(a,b^{\prime\prime})+u(a^{\prime},b)-u(a,b)-u(a^{\prime},b^{\prime\prime})$
and likewise
(7.3) $\int_{a}^{a^{\prime}}\int_{b^{\prime\prime}}^{b^{\prime}}u(s,t)\mathrm{d}t\mathrm{d}s=u(a,b^{\prime})+u(a^{\prime},b^{\prime\prime})-u(a,b^{\prime\prime})-u(a^{\prime},b^{\prime}).$
Adding up these two identities, it follows that (7.1) holds in this instance as well. The assertion of the theorem now follows from the general quadrature identity (7.1). ∎
7.2. Checking the solution formula
We first verify that the purported solution attributed to Riemann actually solves the Darboux-Goursat problem.
Proof of Theorem 1.2.
It is an easy consequence of the formula defining $u=\mathbf{R}[f,g]$ that $u$ gets to be continuous. Moreover, the values along the axis are as required, in accordance with the calculations (1.4) and (1.5).
It remains to check that $u=\mathbf{R}[f,g]$ solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory. To this end, the first step is to realize that
$\partial_{x}\partial_{y}(f(x)+g(y))=0$
holds in the sense of distribution theory. It now follows from the definition of Riemann’s solution that in the sense of distribution theory,
(7.4) $\partial_{x}\partial_{y}u(x,y)=\partial_{x}\partial_{y}(f(x)+g(y))-f(0)\partial_{x}\partial_{y}J_{0,0}(x,y)\\ -\partial_{x}\partial_{y}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s-\partial_{x}\partial_{y}\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t\\ =f(0)J_{0,0}(x,y)-\partial_{x}\partial_{y}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s-\partial_{x}\partial_{y}\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t,$
where we applied the fact that the function $J_{0,0}$ solves the Klein-Gordon equation itself. Next, by Calculus, we have that
(7.5) $\partial_{x}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s=J_{1,0}(y,0)f(x)+\int_{0}^{x}\partial_{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s\\ =f(x)y+\int_{0}^{x}\partial_{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s,$
and, consequently,
(7.6) $\partial_{x}\partial_{y}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s=\\ =f(x)+\int_{0}^{x}\partial_{x}\partial_{y}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s=f(x)-\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s.$
The analogous calculation switching the roles of $x$ and $y$ gives that
(7.7) $\partial_{x}\partial_{y}\int_{0}^{y}J_{1,0}(x,y-t)g(s)\,\mathrm{d}t=g(y)-\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t.$
It now follows by a combination of (7.4), (7.6), an (7.7) that
(7.8) $\partial_{x}\partial_{y}u(x,y)=f(0)J_{0,0}(x,y)-f(x)-g(y)+\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s\\ +\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t=-u(x,y).$
This means that $u=\mathbf{R}[f,g]$ solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory, which completes the proof. ∎
7.3. Obtaining the uniqueness of Riemann’s solution
We now supply the proof of Theorem 1.3, based on the classical Picard method.
Proof of Theorem 1.3.
By considering the difference $v=u-\mathbf{R}[f,g]$, we have a continuous solution to the Klein-Gordon equation $v^{\prime\prime}_{xy}+v=0$ with $v(x,0)=0$ and $v(0,y)=0$. By the quadrature identity (1.7) with $v$ in place of $u$,
(7.9) $v(x,y)=-\int_{0}^{x}\int_{0}^{y}v(s,t)\,\mathrm{d}t\mathrm{d}s.$
Let $Q$ denote the rectangle $a\leq x\leq a^{\prime}$, $b\leq y\leq b^{\prime}$, where the parameters are chosen such that the origin $(0,0)$ is in $Q$. In view of (7.11), we have
(7.10) $|v(x,y)|\leq|xy|\,\|v\|_{L^{\infty}(Q)},\qquad(x,y)\in Q,$
so that if the rectangle $Q$ is not too big, more precisely,
$A:=\max\{|a|,a^{\prime}\}\,\max\{|b|,b^{\prime}\}<1,$
we find that
$\|v\|_{L^{\infty}(Q)}\leq A\|v\|_{L^{\infty}(Q)},$
which forces $\|v\|_{L^{\infty}(Q)}=0$, that is, $v=0$ holds on $Q$. But then we can do a horizontal translation $(x,y)\mapsto(x+\beta,y)$ (for a small $\beta$, positive or negative) and apply the same argument again after the translation. Iteratively, we find that $v=0$ holds in the whole strip $\infty<x<+\infty$, $b\leq y\leq b^{\prime}$. After that we apply instead vertical translations and obtain that $v=0$ throughout. ∎
Proof of Theorem 1.4.
As in the proof of Theorem 1.3 , we have, by the quadrature identity (1.7),
(7.11) $u(x,y)=-\int_{0}^{x}\int_{0}^{y}u(s,t)\,\mathrm{d}t\mathrm{d}s,\qquad a\leq x\leq a^{\prime},\,\,b\leq y\leq b^{\prime},$
which puts us in the same setting as in the proof of Theorem 1.3. The conclusion that $u=0$ on $a,a^{\prime}]\times[b,b^{\prime}]$ follows in an analogous fashion. ∎
Acknowledgements
The second author acknowledges support from Vetenskapsrådet (VR grant 2020-03733).